OpenStudy (anonymous):

Given (n log n)= 10^6 find the value for n

OpenStudy (jiteshmeghwal9):

\[\LARGE{a \log b=\log b^a}\]\[\LARGE{\log x=a \implies 10^a=x}\]use these two properties